# A case study of gas impacted by black-hole jets with the JWST: outflows, bow shocks, and high excitation of the gas in the galaxy IC5063

### v1

@@ 3.2 Regions influenced by the jet and outflows
Bow-shock-shaped structures are revealed for the first time for this galaxy, already for velocities as low as those of the gaseous disk, i.e., 260 km s-1 (as deduced from the S(1) first moment map of Fig. 4; see also Fonseca-Faria et al. 2023). A particularly prominent such structure is the -bright region marked as b1 in Fig. 5, 1200 pc south-east of the nucleus. Even though in the low-resolution S(1) data it appears like a clump, in line with higher resolution data, i.e. in S(3), S(4), and S(5), its shape is very clearly indicative of a bow shock that is headed outwards in the direction of the SE branch of the ionized gas EELR (marked as f3 in Fig. 5). This structure is indeed a bow shock, not a projection effect of spatially unrelated clumps, as it maintains its shape in cubes in which there is no detectable emission by any other structure in its vicinity (such as a weak spiral arm). Another bow shock is b2, with a similar orientation according to S(3), S(4), and S(5). In S(7), a third bow-shock-shaped structure is also seen at the SE lobe itself, at 150 km s-1 (although this structure may be part of the ionized gas bubble b3; see below). The larger-scale structure connecting b1, b2, and b3, as best seen for S(1) in the velocity range 100V400 km s-1, does not follow the spiral patterns described above (marked as sp in Fig. 5), but the jet axis and the EELR edge instead. These findings suggest that the jet has traveled in the EELR and driven bow shocks in regions b1 and b2. In b1 and all along the f3 filament leading up to b1 (shown e.g., in the [Ne III] image at V300 km s-1; Fig. 5), a weak outflow is seen in many emission lines (Fig. 7). In the north-west part of the galaxy, both the edges (f5, f6, f8; Fig. 6) and the interior of the ionized gas EELR (f7 and f9) are rather filamentary. One of these structures, f9, is in a nearly straight line with the jet axis, as seen at 17.8 GHz, again suggesting that the jet has propagated in the EELR.
@@
In the 17.8 GHz radio image presented by Morganti et al. (2007), bright radio emission exists mainly at the radio core, lobes, and between them. No significant emission is detected near b1. In the other side of the galaxy, contours shown close to the NW radio lobe are elongated in a direction comparable to that of the f5 branch outer part, as delineated by [Mg V]. A similar elongation is also seen in contours of a 8 GHz image (Morganti et al., 1998). Therefore, it is possible that some weak radio emission exists, to be revealed with higher sensitivity observations. It is also possible that a jet passed through the region with relativistic speed and that we now see gas pushed and excited by the slower-moving outflows that it raised. As shown in studies of other nearby galaxies, it is possible to detect the passage of a past or of a weak radio jet by the impact it leaves on the ISM: Aalto et al. (2016, 2020), Fernández-Ontiveros et al. (2020), Pereira-Santaella et al. (2022), and Papachristou et al. (2023) demonstrated that outflows can reveal the existence of jets, even if these jets are not seen in radio wavelengths. As we will show below for IC5063, outflows exist at the nucleus, at the radio lobes and in many other regions.
@@
An examination of the cube slices at different velocity ranges indicates that there are, in fact, several starting points of outflows within 1 from each radio lobe. In the overall MIRI FOV, twice as many spatially discrete outflows as those already known from the VLT SINFONI data are now revealed (Dasyra et al., 2015). In the western part of the galaxy, the fastest outflow with positive velocities (of 1000 km s-1) is seen in the clump c2, which is about midway between the nucleus and the NW radio lobe, and which lies at the center of an overpressurized region in the MUSE data (Dasyra et al., 2022). The second highest velocities are found in the clump c1, which is 100(10) pc away from the bifurcation point at the EELR base. At negative velocities, of about -1000 km s-1, yet another clump (marked as c7 in Fig. 6) is seen in [S IV] 90 pc north east of the same bifurcation point, and south of a region with high ratio of outflowing to regularly rotating gas seen in SINFONI [Fe II] 1.644 data (Dasyra et al., 2015). All these regions could either be discrete starting points of outflows or fragmented clouds in the same outflow. Beyond the bifurcation point, a filamentary outflow is detected along the NW EELR edge (see e.g., [Mg V] at -600V-300 km s-1). Similar findings can be reported for the eastern part of the galaxy. The outflow along the f3 filament in the EELR SE edge extends out to 800 pc from the SE bifurcation point. An individual clump can be identified in it (c4). North of it, closer to the other EELR branch (f4), two more clumps that are moving at high velocities are identified (c5,c6). In c6, which is in the EELR interior and close to the jet axis, the velocities reach 900 km s-1 for, e.g., [S IV]. Another outflowing clump closer to the nucleus is c3, seen up to 500 km s-1 for the same line.
@@
The moment maps also indicate the existence of another clump moving at opposite velocities compared to those of the disk. It is a cloud 300 pc north of the NW EELR base, seen in both the first and the second moment maps of [Ar II], [Ne II], and [Fe II]. It is also seen in the dispersion map of the 1.644 [Fe II] line observed with SINFONI. We call it c8 in Fig. 9, where a summary of all jet-cloud interactions is schematically shown.
@@
In total, we see more than ten outflowing clouds and structures in our data. The exact number depends on whether some clumps can be considered part of a common outflow (that started from the same point and that is headed in the same directions) or not, as the medium that we see is very clumpy. A series of jet-cloud interactions and reformation of clouds in outflows (Zubovas & King 2014) could be contributing to the clumpiness of the medium. We note that it is possible that even more outflows exist in our data, but that the cruciform artifact (see Section A.1) prevents us from identifying them with certainty. Such an example is the structure u1. Similarly, some low S/N regions seen around the nuclear PSF in the -1150V-850 km s-1 [S IV] image could either be real or due to PSF residuals.
@@
A plot visualizing all regions with high gas turbulence can be found in Fig. 10. It shows how wide-spread the turbulence in the center of IC5063 is. In total, the gas kinematics are disturbed in 2/3rds of the ch1 FOV. Moreover, turbulent medium appears to be moving in projection towards a dusty structure seen in the optical image of Figs. 1 and 10 that runs nearly parallel to the jet axis, 1200 pc north of it. This dustry structure comprises clouds with clear signs of either ram pressure or radiation pressure having acted on them. It is thus possible that these clouds have been affected by a past episode of jet or AGN radiation feedback.
@@
Another diagram showing how ubiquitous outflows are in our data is the position-velocity diagram along the major axis of rotation, constructed for the bright [S IV] line (Fig. 11). Theoretical curves of the rotational and escape velocities as a function of radius are also overplotted. The gravitational potential associated with the matter distribution in IC5063, at the radii under examination, is assumed to be well approximated by a Hernquist potential (Mukherjee et al., 2018), for which analytic formulas describe the radial profiles of the rotational velocity and the escape velocity (Hernquist, 1990). For the mass and the scale length of the stellar distribution needed for this calculation, we also adopted the parameter values of Mukherjee et al. (2018). These relied on observations of the central stellar velocity dispersion () and effective radius () from Kulkarni et al. (1998) and Morganti et al. (2015), and indicated a stellar mass of and a scale length of for the stellar bulge. Correcting for the inclination of the galaxy, with an inclination angle of (Morganti et al., 2015), we obtained the rotational velocity and escape velocity curves in the position-velocity diagram (Fig. 11). The position angle along which the [S IV] flux was collapsed was identical to that of the major axis of rotation of S(1), 108∘ (Fig. 4). Unlike the molecular gas traced by S(1), which never reaches projected velocities greater than the escape velocity, the outflowing ionized gas probed by [S IV] indeed travels faster than the escape velocity nearly all along the inner 600 parsec of the galaxy. Only in the 100 central parsecs this claim cannot be reliably made due to PSF residuals and the to contribution of the black hole of 2.8108 (Nicastro et al. 2003) to the gravitational potential. In the south-east of the galaxy, the faint outflow seen in Fig. 7 for the b1 bow shock region is also continuous, appearing all along the EELR edge from 600 to 1500 pc from the center (see also Wagner & Appenzeller 1989 for a blue wing in [O III] 5007 Å at 600 pc). However, the gas projected velocities there do not reach the corresponding escape velocities.
@@ 3.3 Density of the ionized gas
The density of the ionized gas can be determined in the MIRI wavelength range with two sets of lines of the same species and ionization potential: the [Ar V] 13.1/7.9 flux ratio and the [Ne V] 24.3/14.3 flux ratio. We use the former ratio, which has sufficient spatial resolution for the creation of a map. The density range that [Ar V] 13.1/7.9 can measure is shown in Fig. 12. Its values are computed with the code Pyneb (Luridiana et al., 2015) for different temperatures and for line intensities that are reflecting electronic levels that are populated based on the Boltzmann distribution, i.e., for gas under LTE. The derived density map (Fig. 12) was computed for an indicative temperature of 10000K. This temperature was deemed appropriate given that flux ratios 1.5 were found in several pixels of our data, and also given that previous temperature calculations from [N II] lines in MUSE data were in the range 7000-12000K (Dasyra et al., 2022). The density map shows discrete regions of overdense ionized gas, with densities somewhat higher than 10000 atoms per cubic centimeter in locations where outflow starting points are seen, including c1 and c2. For comparison, for the vicinity of c2, Holden et al. (2023) report densities of 6000 cm-3 based on [Ar IV] in the optical. Similarly, in the NW EELR (regions f6, f7), in a region with outflows in our data and with particular polarization in HST FOC data (potentially due to the outflows; Barnouin et al. 2023), the density seems elevated. Overall, it is noteworthy that the [Ar V] diagnostic ratio, which is applicable at rather high density values (1000 cm-3; Fig. 12) and ionization potential (59.8 eV), works well for so many regions of the nucleus of IC5063. This result could plausibly be indicative of compression of the ionized medium.
@@ 4.1 Ionized gas
In the previous Section, we demonstrated that the jet plays a major role in the observed gas properties in the center of IC5063. A series of outflows along and perpendicular to the known jet axis from radio wavelengths cause great turmoil in the gas kinematics. Bow shocks out to 1200 pc and filamentary outflows all along the ionized gas EELR SE branch, as well as outflows out to 800 pc along the ionized gas EELR NW branch, suggest that the bifurcating appearance of the EELR in both sides of the galaxy is due to jet scattering at the radio lobes. Our findings corroborate that the jet is far more important than the AGN radiation in exciting the gas than previously thought (Morganti et al., 2007; Mingozzi et al., 2019; Venturi et al., 2021). Of course, the AGN radiation may still be contributing to the gas excitation in regions previously cleared out by the jet, as indicated by the existence of AGN radiation and jet inflated cavities in X-ray data (Travascio et al., 2011) and the existence of circumnuclear high-ionization line emission in our MIRI data. Still, the jet contribution to the gas kinematics and excitation is now unambiguously proven even further away than the radio lobes. Already since the presentation of the EELR of IC5063, Colina et al. (1991) pointed out that the radial profile of the ionized gas emission does not agree with the radial profile of the ionized gas density illuminated by a nuclear point source (i.e., the AGN). Instead, these two radial profiles require the existence of an extended ionization source, i.e., of a radio jet, in order to agree. Given that in the eastern part of IC5063, the SE branch is the one with the highest luminosity and most extended outflow of the two EELR edges, and that, in the western part of IC5063, the same holds for the NW branch, it is possible that the jet propagated at a higher position angle than that of the disk. This scenario is also supported by the ionized gas excitation, if examined with the aid of the flux ratio of two bright lines, [Ne III] and [Ne II] (shown in Fig. 13). A similar result was seen in the optical for the [O III]/Hbeta flux ratio (Fonseca-Faria et al., 2023).
@@ 4.2 The molecular gas excitation
In the center, i.e., in the region with detectable radio emission, the gas excitation can be elevated even in the case of shocks alone, because of superposition. Several shocks may be traveling in the ISM at different directions and velocities thanks to the jet scattering at various locations and thanks to the propagation of a slower moving, mass loaded gas cocoon that is linked to the heavier outflow (Morganti et al., 2015). If shocks significantly contribute to the heating, the gas should also be more turbulent, showing broader line profiles in regions where the gas is hotter. Indeed, this is what we observe by comparison with the velocity dispersion maps of Fig. 8.
@@
Both in the vicinity of the radio lobes and in the biconical outflow there is a flux ratio of the total emission (in all lines) over the total PAH emission (in the two detected complexes) that is elevated by a factor of 5-10 compared to regions further out, in the EELR. Even though this result could be reflecting a potential destruction of the PAHs by the jets, and even though the PAH emission is rather weak in the spectrum of Fig. 2, such an interpretation is contradicted by the finding that the PAH emission is bright along the jet axis (see also Diamond-Stanic & Rieke 2010; Ogle et al. 2024). The PAH properties and excitation will too be discussed in a future paper, in the light of the gas excitation.

### v2

@@ 3.2 Regions influenced by the jet and outflows
Bow-shock-shaped structures are revealed for the first time for this galaxy, already for velocities as low as those of the gaseous disk, i.e., 260 km s-1 (as deduced from the S(1) first moment map of Fig. 3; see also Fonseca-Faria et al. 2023). A particularly prominent such structure is the -bright region marked as b1 in Fig. 4, 1200 pc south-east of the nucleus. Even though in the low-resolution S(1) data it appears like a blob (Fig. 3), in lines with higher resolution data, i.e. in S(3), S(4), and S(5), its shape is very clearly indicative of a bow shock (b1 in Fig. 4) that is headed outwards in the direction of the SE branch of the ionized gas EELR (marked as f3 in Fig. 4). This structure is indeed a bow shock, not a projection effect of spatially unrelated clumps. The bow shock maintains its shape in all of S(3), S(4), and S(5), i.e., in cubes in which there is no detectable emission by any other similar-velocity structure in its vicinity: the disk emission, including some weak spiral arm emission, has already faded by S(3). No associated stellar component can be seen in optical and near infrared data at the location of b1: this is a gaseous structure seen in only. Another bow shock is b2, with a similar orientation according to S(3), S(4), and S(5). In S(7), a third bow-shock-shaped structure is also seen at the SE lobe itself, at 150 km s-1 (although this structure may be part of the ionized gas bubble b3; see below). The larger-scale structure connecting b1, b2, and b3, as best seen for S(1) in the velocity range 100V400 km s-1, does not follow the spiral patterns described above (marked as sp in Fig. 4), but the jet axis and the EELR edge instead. These findings suggest that the jet has traveled in the EELR and driven bow shocks in regions b1 and b2. In b1 and all along the f3 filament leading up to b1 (shown e.g., in the [Ne III] image at V300 km s-1; Fig. 4), a weak outflow is seen in many emission lines (Fig. 6). In the north-west part of the galaxy, both the edges (f5, f6, f8; Fig. 5) and the interior of the ionized gas EELR (f7 and f9) are rather filamentary. One of these structures, f9, is in a nearly straight line with the jet axis, as seen at 17.8 GHz, again suggesting that the jet has propagated in the EELR.
@@
In the 17.8 GHz radio image presented by Morganti et al. (2007), bright radio emission exists mainly at the radio core, lobes, and between them. No significant emission is detected near b1. In the other side of the galaxy, contours shown close to the NW radio lobe are elongated in a direction comparable to that of the outer part of the f5 filament, at a PA of 10∘, as it is seen in [Mg V] (Fig. 5) . A similar elongation is also seen in contours of a 8 GHz image (Morganti et al., 1998). Therefore, it is possible that some weak radio emission exists, to be revealed with higher sensitivity observations. It is also possible that a jet passed through the ISM, in the past, with a relativistic speed, and that we now see gas pushed and excited by the slower-moving outflows that it raised. As shown in studies of other nearby galaxies, it is possible to detect the passage of a past or of a weak radio jet by the impact it leaves on the ISM: Aalto et al. (2016, 2020), Fernández-Ontiveros et al. (2020), Pereira-Santaella et al. (2022), and Papachristou et al. (2023) demonstrated that outflows can reveal the existence of jets, even if these jets are not seen in radio wavelengths. As we will show below for IC5063, outflows exist at the nucleus, at the radio lobes and in many other regions, including the b1 bow shock.
@@
An examination of the cube slices at different velocity ranges indicates that there are, in fact, several starting points of outflows. In the overall MIRI FOV, twice as many spatially discrete outflows as those already known from the VLT SINFONI data are now revealed (Dasyra et al., 2015). In the western part of the galaxy, the fastest outflow with positive velocities (of 1000 km s-1) is seen in the clump c2 (Fig. 4), which is about midway between the nucleus and the NW radio lobe, and which lies at the center of an overpressurized region in the MUSE data (Dasyra et al., 2022). The second highest velocities are found in the clump c1, which is 100(10) pc away from the bifurcation point at the EELR base. At negative velocities, of about -1000 km s-1, yet another clump (marked as c7 in Fig. 5) is seen in [S IV] 90 pc north east of the same bifurcation point, and south of a region with high ratio of outflowing to regularly rotating gas seen in SINFONI [Fe II] 1.644 data (Dasyra et al., 2015). All these regions could either be discrete starting points of outflows or fragmented clouds in the same outflow. Beyond the bifurcation point, a filamentary outflow is detected along the NW EELR edge (see e.g., [S IV] at -850V-300 km s-1 and [Mg V] at -600V-300 km s-1 in Fig. 5). Similar findings can be reported for the eastern part of the galaxy. The outflow along the f3 filament in the EELR SE edge extends out to 800 pc from the SE bifurcation point (see [Ne III] at 300V500 km s-1 in Fig. 4). An individual clump can be identified in it (c4). In two more high velocity clumps, one close to the f4 edge of the SE EELR (clump c5), and another close to the jet axis (clump c6), the velocities reach 900 km s-1 for [S IV]. Another outflowing clump closer to the nucleus is c3, seen up to 500 km s-1 for the same line.
@@
The moment maps also indicate the existence of another clump moving at opposite velocities compared to those of the disk. It is a cloud 300 pc north of the NW EELR base, seen in both the first and the second moment maps of [Ar II], [Ne II], and [Fe II]. It is also seen in the dispersion map of the 1.644 [Fe II] line observed with SINFONI. We call it c8 in Fig. 7 (see the [Ar II] panel).
@@
In total, we see more than ten outflowing clouds and structures in our data. The exact number depends on whether some clumps can be considered part of a common outflow (that started from the same point and that is headed in the same directions) or not, as the medium that we see is very clumpy. A series of jet-cloud interactions and reformation of clouds in outflows (Zubovas & King 2014) could be contributing to the clumpiness of the medium. We note that it is possible that even more outflows exist in our data, but that the cruciform artifact (see Section B.1) prevents us from identifying them with certainty. Such an example is the structure u1 (Fig. 4). Similarly, some low S/N regions seen around the nuclear PSF in the -1150V-850 km s-1 [S IV] image could either be real or due to PSF residuals.
@@
A plot visualizing all regions with high gas turbulence can be found in Fig. 9. It shows how wide-spread the turbulence in the center of IC5063 is. In total, the gas kinematics are disturbed a region of 2.21.6 kpc2. Moreover, turbulent medium appears to be moving in projection towards a dusty structure seen in the optical image of Figs. 1 and 9 that runs nearly parallel to the jet axis, 1200 pc north of it. This dusty structure is seen as a wavy interface and pillar-shaped clouds, which could be associated with radiation pressure or ram pressure (and associated instabilities, such as the Vishniak instability). Winds travelling along the biconical outfow, or along the axis indicated by the [Ne V] and [S IV] velocity dispersion contour elongation (i.e., at a 15∘ position angle from the NW radio lobe), could have acted on this structure, giving it its current shape. It is thus possible that these clouds have been affected by a past episode of jet or AGN radiation feedback.
@@
Another diagram showing how ubiquitous outflows are in our data is the position-velocity diagram along the major axis of rotation, constructed for the bright [S IV] line (Fig. 10). Theoretical curves of the rotational and escape velocities as a function of radius are also overplotted. The gravitational potential associated with the matter distribution in IC5063, at the radii under examination, is assumed to be well approximated by a Hernquist potential (Mukherjee et al., 2018), for which analytic formulas describe the radial profiles of the rotational velocity and the escape velocity (Hernquist, 1990). For the mass and the scale length of the stellar distribution needed for this calculation, we also adopted the parameter values of Mukherjee et al. (2018). These relied on observations of the central stellar velocity dispersion () and effective radius () from Kulkarni et al. (1998) and Morganti et al. (2015), and indicated a stellar mass of and a scale length of for the stellar bulge. Correcting for the inclination of the galaxy, with an inclination angle of (Morganti et al., 2015), we obtained the rotational velocity and escape velocity curves in the position-velocity diagram (Fig. 10). The position angle along which the [S IV] flux was collapsed was identical to that of the major axis of rotation of S(1), 108∘ (Fig. 3). Unlike the molecular gas traced by S(1), which never reaches projected velocities greater than the escape velocity, the outflowing ionized gas probed by [S IV] indeed travels faster than the escape velocity nearly all along the inner 600 parsec of the galaxy. Only in the 100 central parsecs this claim cannot be reliably made due to PSF residuals and the to contribution of the black hole of 2.8108 (Nicastro et al. 2003) to the gravitational potential. In the south-east of the galaxy, outflowing gas is continuously distributed from 600 to 1500 pc from the nucleus, appearing all along the EELR edge. It can be seen in the position-velocity diagram and in Fig. 9 with the aid of the [Ne III] contours leading up to the b1 bow shock. Weak line wings associated with this nearly-kiloparsec-long outflow can be seen in Fig. 6, near b1. There, the projected gas velocities do not reach the corresponding escape velocities. At 600 pc, the other end of this outflow, a blue wing in [O III] 5007 Å has long been known to exist (Wagner & Appenzeller, 1989). At distances below 600 pc, the filament f10 seen for S(5) (Fig. 5), which is nearly parallel to the jet axis, likely contains outflowing gas.
@@ 3.3 Density of the ionized gas
The density of the ionized gas can be determined in the MIRI wavelength range with two sets of lines of the same species and ionization potential: the [Ar V] 13.1/7.9 flux ratio and the [Ne V] 24.3/14.3 flux ratio. We use the former ratio, which has sufficient spatial resolution for the creation of a map. The density range that [Ar V] 13.1/7.9 can measure is shown in Fig. 11. Its values are computed with the code Pyneb (Luridiana et al., 2015) for different temperatures and for line intensities that are reflecting electronic levels that are populated based on the Boltzmann distribution, i.e., for gas under LTE. The derived density map (Fig. 11) was computed for an indicative temperature of 10000K. This temperature was deemed appropriate given that flux ratios 1.5 were found in several pixels of our data, and also given that previous temperature calculations from [N II] lines in MUSE data were in the range 7000-12000K (Dasyra et al., 2022). The density map shows discrete regions of overdense ionized gas, with densities somewhat higher than 10000 atoms per cubic centimeter in locations where outflow starting points are seen, including c2. For comparison, for the vicinity of c2, Holden et al. (2023) report densities of 6000 cm-3 based on [Ar IV] in the optical. Similarly, in the NW EELR (regions f6, f7), in a region with outflows in our data and with particular polarization in HST FOC data (potentially due to the outflows; Barnouin et al. 2023), the density seems elevated. Overall, it is noteworthy that the [Ar V] diagnostic ratio, which is applicable at rather high density values (1000 cm-3; Fig. 11) and ionization potential (59.8 eV), works well for so many regions of the nucleus of IC5063. This result could plausibly be indicative of compression of the ionized medium.
@@ 4.1 Ionized gas
In the previous Section, we demonstrated that the jet plays a major role in the observed gas properties in the center of IC5063. A series of outflows along and perpendicular to the known jet axis from radio wavelengths cause great turmoil in the gas kinematics. Bow shocks out to 1200 pc and filamentary outflows all along the ionized gas EELR SE branch, as well as outflows out to 800 pc along the ionized gas EELR NW branch, suggest that the bifurcating appearance of the EELR in both sides of the galaxy is due to jet scattering at the radio lobes. Our findings indicate that the jet is far more important than the AGN radiation in exciting the gas than previously thought (Morganti et al., 2007; Mingozzi et al., 2019; Venturi et al., 2021). Of course, the AGN radiation may still be contributing to the gas excitation in regions previously cleared out by the jet, as indicated by the existence of AGN radiation and jet inflated cavities in X-ray data (Travascio et al., 2011) and the existence of circumnuclear high-ionization line emission in our MIRI data. Still, the jet contribution to the gas kinematics and excitation is now unambiguously proven even further away than the radio lobes. Already since the presentation of the EELR of IC5063, Colina et al. (1991) pointed out that the radial profile of the ionized gas emission does not agree with the radial profile of the ionized gas density illuminated by a nuclear point source (i.e., the AGN). Instead, these two radial profiles require the existence of an extended ionization source, i.e., of a radio jet, in order to agree. It is even possible that the jet has traveled at a greater position angle than that of the disk (i.e., at an orientation more inclined towards NW-SE), because the gas emission is brightest in the SE EELR branch east of the nucleus and in the NW EELR branch west of the nucleus. The ionized gas excitation, when examined with the aid of the flux ratio of two bright lines, [Ne III] and [Ne II] (shown in Fig. 12), shows that these two branches also contain the most excited gas (see also a similar result for the [O III]/Hbeta ratio in the optical; Fonseca-Faria et al. 2023). Lines of high IP, such as [Fe VIII] (with 99.1eV) emit along the same EELR branches in our Webb data, and outflows are detected there too.
@@ 4.2 The molecular gas excitation
In the center, i.e., in the region with detectable radio emission, the gas excitation can be elevated even in the case of shocks alone, because of superposition. Several shocks may be traveling in the ISM at different directions and velocities thanks to the jet scattering at various locations and thanks to the propagation of a slower moving, mass loaded gas cocoon that is linked to the heavier outflow (Morganti et al., 2015). If shocks significantly contribute to the heating, the gas is also likely to be more turbulent, showing broader line profiles in regions where the gas is hotter. The comparison to the velocity dispersion maps of several bright lines in Fig. 7 advocates in favour of this scenario.
@@
The contribution of lines from non-fully-thermalized states is instructive when studying the -to-PAH flux ratio. A map of this ratio, when constructed from the thermal emission of the S(1), S(2), and S(3) lines (Fig. 14; left), reveals that the highest value is seen at the b1 bow shock. The elevated thermal over PAH emission found in numerous radio galaxies (Ogle et al., 2010) can, thus, indeed be attributed to shocks. A map of the same ratio, when constructed from the total (thermal and fluorescent) emission of all lines (Fig. 14; right) is instead brightest at c1, close to the north radio lobe. There, the maximum detected value (3.5) is 10 times greater than the average EELR value. This result could reflect the higher excitation of the molecular gas at the vicinity of the radio lobes, as presented above. Alternatively, an elevated -to-PAH ratio could signify the destruction of PAHs (e.g., Diamond-Stanic & Rieke 2010), even though the PAH emission is bright along the jet axis (see also Ogle et al. 2024). As the -to-PAH ratio is an output of shock models (e.g., Kristensen et al. 2023; Villa-Velez et al. 2024), the PAH properties will be further discussed in our forthcoming paper.